A friend of yours is planning to open a unique Mexican-Lebanese fusion food restaurant in Barrie, Ontario (population 150,000). 


Given the restaurant's unique fusion food options, along with the popularity of regular Mexican and Lebanese food cuisines, your friend expects to have a slight advantage over most of the other 240 restaurants in the city, even though all restaurants have the same geographic access to the city's populace.The Canadian restaurant and food service association estimates that total industry revenues for Canada (population 37.4 million) will be $75 billion next year.


What would be the Total Addressable Market (TAM), Serviceable Available Market (SAM) and Serviceable Obtainable Market (SOM) for this restaurant?


1
Expert's answer
2021-10-25T17:53:17-0400

Solution:

The total addressable market (TAM) for Canada is the total revenue generated by all industries.

It is the total available market, or the total revenue opportunity for a product or service if it achieves 100% market share.

The Total Addressable Market (TAM) is estimated to be worth $75 billion.

 

Service Available Market (SAM) = Revenue per person "\times" Barrie, Ontario Population

Revenue per person = "\frac{7,500}{37.4}" = 200.54

Barrie, Ontario population = 150,000

Service Available Market (SAM) = 200.54 "\times"150,000 = $30.081 million

 

Serviceable Obtainable Market (SOM) = Service Available Market (SAM) "\div" No. of competitors

No. of competitors = 200

Serviceable Obtainable Market (SOM) = "\frac{30.081}{200}" = $150,405